OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.9 megabytes (MB). The size of the high-quality version is 4.5 MB. Yesterday, there were 1210 downloads of the song, for a total download size of 4869 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

\[ 2.9x+4.5y=4869\\ x+y=1210 \]

OpenStudy (anonymous):

Where \(x\) is standard version sales and \(y\) is hq version sales.

OpenStudy (anonymous):

@wio okay

OpenStudy (anonymous):

multiply the bottom equation by \(-4.5\). What do you get?

OpenStudy (anonymous):

The x+y=1210 one? @wio

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

-5445 @wio

OpenStudy (anonymous):

What is the entire equation though?

OpenStudy (anonymous):

-4.5x-4.57=-5445? @wio

OpenStudy (anonymous):

Okay, now add that equation to the top equation.

OpenStudy (anonymous):

1.6x=576 @wio

OpenStudy (anonymous):

Okay what is next?

OpenStudy (anonymous):

x=360

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

that's my answer?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

@wio